Find the 12th term of the sequence an = n(n - 6).
Answer by stanbon(75887)

Find the 12th term of the sequence an = n(n - 6)
-------
a(12) = 12(12-6) = 12*6 = 72
--------------------------------
Cheers,
Stan H.